Shift report 20011229 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Sat Dec 29 2001 - 08:42:29 EST

  • Next message: Flemming Videbaek: "unfilled shifts Brahms"

    ------------------------------------------------------------
    Date: Sat Dec 29 08:42:29 2001, Shift: 00:00-08:00
    Supervisor: hiro, with: 
    Log book#11 pages 111 - 111
    
    Shift Summary:
    ------------------------------------------------------------
    1.  Beam is stored until 1:00
        Beam is stored at 5:40. 
        (It is still going.)
    2.  Run 6190, 6192, 6193
    3. None detected.
    
    For silicon:
     To test more about attenuation for 
    silicon, the 50 ohm resister (terminator) in
    Si1_1c is removed(It is still back terminted.)
    The attenuation is 1.5x (not quite 1x)
    
    
    
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Sat Dec 29 2001 - 08:44:34 EST